OpenStudy (anonymous):

one number is 4 times as large as another.When the smaller is decreased by 2 it is 14 less than the larger number. what are the 2 numbers?

OpenStudy (radar):

x=smaller number 4x=larger number x-2 =smaller number decreased by 2 4x-14=larger number less 14 x-2=4x-14 given in the problem info. x-4x=-14+2 -3x=-12 x=4 smaller number 4x=16 larger number